OpenStudy (anonymous):

ten litres of 20% acid solution are mixed with 30 litres of 30% of acid solution. which of the following is the percent of acid in the final mixture? a) 21% b)22.5% c)25% d)27.5% e)28%

OpenStudy (anonymous):

Equation: acid + acid = acid 0.20*10 + 0.30*30 = x(40) --- Multiply thru by 100 to get: 20*10 + 30*30 = 4000x 200 + 900 = 4000x x = 1100/4000 x = 0.275 or 27.5%

OpenStudy (whpalmer4):

10 liters of 20% acid solution contains how many liters of acid? 30 liters of 30% acid solution contains how many liters of acid? add those two numbers together, and divide by the total number of liters of acid solution in the final mixture to get your percentage.

OpenStudy (anonymous):

So D

OpenStudy (anonymous):

>.> You were checking my equation to make sure it was right

OpenStudy (whpalmer4):

@DragonFlame nah, I was typing at the same time you were typing, apparently

OpenStudy (anonymous):

Oh i c